Sal's Rare Roses is a single price monopoly. The table shows the demand schedule for Sal's Rare Roses (columns 1 and 2) and the firm's total cost schedule (columns 2 and 3) What is Sal's maximizing output, price, and economic profit?

Price (dollars per bush) Quantity (bushes per hour) Total Cost (dollars per hour)
12 0 1
11 1 6
10 2 13
9 3 22
8 4 33
7 5 46
Answer #1
P Q TC TR MC MR
12 0 1 0
11 1 6 11 5 11
10 2 13 20 7 9
9 3 22 27 9 7
8 4 33 32 11 5
7 5 46 35 13 3

TR=P*Q

MC=Change in TC/change in Q

MR=Change in TR/change in Q

The monopoly firm will set MC=MR for profit maximization

Output = 2 units

Price = 10

Profits = TR-TC = 20-13 = 7